OpenStudy (anonymous):

Can someone please help me simplify this integral for a Calc3 class. Im supposed to simplify then look it up on table of integrals. Integral Cosx/Sin^2x=Sinx dx

OpenStudy (anonymous):

\[ \int \frac{\cos x }{\sin^2 x}dx = \int \cot(x)\csc(x)\;dx=\int -d[\csc x] \]

OpenStudy (anonymous):

Since \(d[\csc x]=-\cot x \csc x\;dx\)

OpenStudy (anonymous):

Oh wait, something tells me that isn't supposed to be an equals sign... please right the integrand properly.

OpenStudy (anonymous):

sorry I typed that wrong it was supposed to be \[\int\limits \cos x/ \sin^2 x + \sin x\]

OpenStudy (anonymous):

So this equals \[\int\limits \cot x / (sinx)+1\]

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

Then that becomes \[\int\limits \cot x \csc (x+1) dx\] ?